What is the average power output of a weightlifter who can lift 220 kg to a height of 1.7 m in 1.9 s?
1
Expert's answer
2019-02-07T12:15:28-0500

By the definition of the power we have:

"P = \frac{W}{t} = \frac{Fh}{t} = \frac{mgh}{t} = \frac{220 kg \cdot 9.8 \frac{m}{s^2} \cdot 1.7 m}{1.9 s} = 1929 W."

Answer:

"P = 1929 W"

.